Two-point charges are arranged on the x-y coordinate system as follows; q1= 3.0 nC at (0, 3m) and q2= -9.0 nC at (0,0). Find the electric field at point on x-axis 4.0 m from the origin. 


1
Expert's answer
2022-03-21T14:06:45-0400

The net electric field

"\vec E=\vec E_1+\vec E_2"

"\vec E=k\frac{q_1}{r_1^3}\vec r_1+k\frac{q_2}{r_2^3}\vec r_2"

"\vec r_1=(4,-3),\quad r_1=\sqrt{4^2+(-3)^2}=5"

"\vec r_2=(4,0),\quad r_2=\sqrt{4^2+0^2}=4"


"\vec E=9\times 10^9\times \frac{3.0\times 10^{-9}}{5^3}(4,-3)\\
+9\times 10^9\times \frac{-9.0\times 10^{-9}}{4^3}(4,0)\\
=(0.864,-0.648)+(-5.063,0)\\
=(-4.2,-0.65)\:\rm N/C"

The magnitud of field

"E=\sqrt{E_x^2+E_y^2}=\sqrt{(-4.2)^2+(-0.65)^2}\\
=4.25\:\rm N/C"